OpenStudy (anonymous):

Find the quotient (22/7) / (11/14)

OpenStudy (igreen):

4

OpenStudy (kainui):

\[\frac{ \frac{ 22 }{ 7 } }{ \frac{ 11 }{ 14 } }\] When you're dividing fractions, you flip the second fraction and multiply.

OpenStudy (anonymous):

preciate it